OpenStudy (anonymous):

how to solve for pi in SA=2*pi*r^2+2*pi*r*h

OpenStudy (anonymous):

\[SA=2\pi r^2 + 2\pi rh\]Factor \(\pi\) out of the right hand side.\[SA=\pi \left( 2r^2+2rh \right)\]Divide both sides by \(2r^2+2rh\)\[\pi = \frac{ SA }{ 2r^2+2rh }\]